# Grasp EveryThing (GET): 1-DoF, 3-Fingered Gripper with Tactile Sensing for Robust Grasping

### v2

@@ Abstract
We introduce the Grasp EveryThing (GET) gripper, a novel 1-DoF, 3-finger design for securely grasping objects of many shapes and sizes. Mounted on a standard parallel jaw actuator, the design features three narrow, tapered fingers arranged in a two-against-one configuration, where the two fingers converge into a V-shape. The GET gripper is more capable of conforming to object geometries and forming secure grasps than traditional designs with two flat fingers. Inspired by the principle of self-similarity, these V-shaped fingers enable secure grasping across a wide range of object sizes. Further to this end, fingers are parametrically designed for convenient resizing and interchangeability across robotic embodiments with a parallel jaw gripper. Additionally, we incorporate a rigid fingernail to enhance small object manipulation. Tactile sensing can be integrated into the standalone finger via an externally-mounted camera. A neural network was trained to estimate normal force from tactile images with an average validation error of 1.3 N across a diverse set of geometries. In grasping 15 objects and performing 3 tasks via teleoperation, the GET fingers consistently outperformed standard flat fingers. Finger designs for use with multiple robotic embodiments are available on GitHub.
[…]
figureGrasp EveryThing (GET) Gripper. (a) Holding a toy hammer with simultaneous tactile image recorded from HDR camera at the back of the single finger side. (b) The GET gripper is designed to be easily integrated onto parallel grippers commonly found on existing in-the-wild data collection systems with added torque constraint. We demonstrate its capability in completing a variety of household tasks through teleoperation on the ALOHA system [1].
Figure 1: Assembly of GET components. Left: Rendered exploded view of mechanical and electronic components in each finger, arranged in order of assembly. Parts that are made of Onyx are 3D printed on Markforged printer. Right: Fabricated set of GET fingers with rigid fingernails highlighted.
@@ I Introduction
By introducing a third finger onto a parallel jaw gripper, we can create a more robust and universally capable gripper that remains just as easy to implement and control. The contact patch provided by this additional finger works to form a lever arm to better resist torques about both normal and frictional directions. With V-shaped fingers, the length of this lever arm can vary as needed in correspondence to the shape or size of grasped objects.
@@ II-A Robotic Gripper Design
Rigid, parallel jaw grippers are easy to model and control due to their 1-DoF. Recently, they have become especially prevalent given modern learning-from-demonstration style control policies that heavily rely on teleoperated demonstration and in-the-wild data collection [5, 2]. However, traditional flat-fingered parallel jaw grippers often form only two contact patches on grasped objects. This limits their ability to constrain against torques and achieve secure grasps, making them less effective for dynamic tasks and tool usage [7, 10]. Alternative finger designs have been developed in attempt to overcome this obstacle. Most notable are compliant structure fingers that passively wrap around objects to increase contact area and further geometrically constrain grasped objects [11, 12, 13]. These grippers are favorable due to their improved performance over rigid fingers but still often fail in dynamic tasks requiring strong torque resistance. Other form factors have been developed to increase torque resistance beyond the traditional 2-fingered parallel gripper [1, 14].
By introducing additional DoF, it is possible to design grippers that are more dexterous, anthropomorphic, and capable of handling more dynamic tasks [15, 16, 17, 18]. The increased complexity of these grippers empowers more performative action, including strong multi-contact grips and in-hand manipulation actions. Unfortunately, hand designs often become bulky, costly, and scale to large numbers of DoF that leave them difficult to control. Crucially, hands are complicated to teleoperate [19, 20], which hinders the development of learning-from-demonstration control policies. Naturally, many have sought to minimize DoFs while maximizing dexterous capabilities through underactuation [21]. However, modeling underactuation can be challenging and often teleoperation remains nontrivial.
@@ II-B Camera-Based Tactile Sensors
Recently, researchers have developed approaches where a single camera is used to capture both tactile and proximity data, capturing objects and their associated contact geometry in the same image [28, 29]. This eliminates the need for a wrist-mounted camera. However, an externally-mounted camera may be susceptible to lighting changes, introducing noise that could inhibit the ability to reconstruct surface depth as precisely as GelSight sensors traditionally do [23]. This is often acceptable since many learning-based implementations only use RGB tactile images as input, not 3D depth [27, 9, 28].
@@ III-A1 Secure Grasping
Even when grasping more regularly-shaped objects, traditional parallel jaw grippers with two flat fingers struggle to strongly resist torques. We analyze the forces involved in grasping a hammer with a flat, cylindrical handle in Fig. 2b. Here, we assume fingers are not fully rigid and will form patch contacts with the hammer. The 2-finger gripper is able to withstand torque disturbances about the $\hat{x}$ and $\hat{z}$-axes in proportion to the width of the fingers $w$. When a third finger is introduced, the ability to withstand torque additionally depends on the distance between fingers $L$. By designing a parallel gripper with three fingers, we can achieve greater torque compensation about both $\hat{x}$ and $\hat{z}$-axes without the need for overly wide fingers, which may hinder performance in smaller-scale tasks. Resistance to torques in the $\hat{y}$ direction will remain equivalent for both finger configurations and depend on object width.
With our 3-fingered configuration, it is important to note that the greatest resistance against torque disturbances functions about the $\hat{x}$ axis along the length of the fingers. Resistance against disturbances in about the grasping axis $\hat{z}$ will rely on friction, with maximum allowed torque scaling down by the frictional coefficient $\mu$. Object should be grasped strategically to ensure favorable torque compensation.
Beyond incorporating a third finger to improve geometric conformity and increase torque resistance, we design fingers with soft silicone gel pads to further enhance these capabilities. When these soft fingerpads come in contact with a grasped object, they will deform to the local geometry of the object. This works to increase the area of contact patches and grants flexibility in grasping irregularly-shaped objects. Additionally, depending on the geometry of the object, an indentation formed at the contact patch may provide resistance along the face of the finger, with the elastic silicone acting as a spring to resist object movement.
@@ III-A2 Self-Similarity & Scalability
On the 2-finger side of our gripper, the fingers come together to form a V-shape. This creates a varying distance between fingers $L$ along their length to enable secure grasping of many object sizes, as shown in Fig. 3. Larger $L$ at the base grants greater compensation against torque disturbances, while smaller $L$ at the tip enables grasping of small objects. With this design, it is expected that larger objects are grasped at the base of the fingers and smaller objects at the tip.
@@
To prevent fingers from bending outward and diverging, which could lead to grasp instability, we pitch them inward by angle $\theta$. Importantly, fingers are designed to interdigitate. This implies they can freely bend flat to a $0^{\circ}$ pitch angle through contact without interfering with each other, allowing them to grasp even infinitesimally thin objects along their length.
Fingers are parametrically designed such that they can be resized for compatability with any robotic arm. This is done by adjusting a finger length dimension and modifying the attachment geometry. Throughout this paper, we refer specifically to a set of fingers developed for the ALOHA teleoperation system [1]. But, to demonstrate the scalability of this design, additional fingers were fabricated for use with a Franka Emika Panda arm, as shown in Fig. 5.
@@ III-B Optical Design
Camera-based tactile sensing is integrated into the single finger of the gripper. Soldered arrays of RGB LED are placed along the edges of the finger to illuminate the sensor pad. Our finger is significantly longer than it is wide. To address this, lensed blue LEDs are used at the base of the finger to direct light further along its length. In contrast, red and green SMD LEDs are placed on either side of the finger and equipped with diffusion filters to promote uniform illumination.
As discussed in Section III-A2, the design is meant to be parametrically scalable. In accordance with this, each LED board could be replaced by flexible COB LED strips, which can be customized to different lengths, eliminating the need to design a new PCB board for each new finger size.
@@ IV-A Tactile Sensing
Tactile images were collected alongside corresponding measurements from a force probe over varying indentation geometries. Indentation geometries include 3 simple shapes and 15 common objects, such as a coin, USB stick, Lego brick, and more. The force probe used for this experiment was designed with an embedded MMS101 6-axis force torque sensor. CAD files for this probe are available on GitHub. To collect data, we pressed our single finger tactile sensor into the probe multiple times for each geometry while streaming tactile images and recording force measurements for each frame, as shown in Fig. 6b. Pressing forces ranged from 0 to 30 N over the collected dataset. In total, 17.5k images were collected. For reference, while streaming at 30 Hz, 5k images can be collected in under 3 min.
@@ IV-B Performance in Grasping Objects
For these experiments, we utilized the ALOHA system [1] retrofitted with an enhanced gripper design from the ALOHA 2 [6]. We compared three different finger configurations: our proposed design, ALOHA ViperX fingers, and a custom-designed set based on traditional parallel jaw fingers. The traditional fingers are flat and rectangular with a rigid backing and a soft curved gel pad that mimics GET fingers. The ALOHA ViperX fingers are fully rigid with a triangular design that can provide up to four points of contact. We applied high-friction tape across the entire surface of the ALOHA fingers to maximize their capabilities.
@@ IV-C Performance in Manipulation Tasks
Task performance for each set of fingers is plotted in Fig. 9. Our fingers outperformed baseline fingers, specifically in the Toy Hammer and Peanut Butter Spread tasks. The traditional fingers struggled to resist torque moments in these tasks given only two small contact patches. Furthermore, they could not reliably pick small objects in the Car Part Assembly task. The ALOHA ViperX fingers performed better in the Car Part Assembly task, but struggled in the Peanut Butter Spread task. These rigid fingers could not conform to the butter knife’s convex geometry to securely grasp it, as displayed in Fig. 2a.
The GET fingers more successfully completed all tasks than baseline fingers. In the Car Part Assembly task, their fingernails helped pick small pieces from clutter and place them firmly into the desired location. In the Peanut Butter Spread task, the three-finger design could securely grasp the irregularly-shaped butter knife through contact with the cracker. Fingernails were also helpful in this task for removing the peanut butter lid and picking a cracker from the stack. In the Toy Hammer task, GET fingers outperformed the ALOHA ViperX fingers, even with the hammer’s flat, cylindrical geometry. GET fingers had greater capacity to resist the high torque that is needed to pry the nail from the board, as shown in Fig. 2b. Across all tasks, we observe decreased completion time with the GET fingers, suggesting improved ease of use and greater teleoperator confidence.
@@ V Conclusion
We present the GET gripper, a novel 3-fingered design for parallel jaw actuators to more securely grasp objects of many shapes and sizes. This design has been shown to outperform standard 2-fingered designs, particularly in grasping small objects and securely holding larger ones, such as tools for dynamic tasks. Additionally, high-resolution tactile imaging is integrated with force sensing capabilities. The GET gripper could be deployed with the ALOHA system [1] or other teleoperation interfaces to collect rich data for training policies over a wide range of tasks.

### v3

@@ Abstract
We introduce the Grasp EveryThing (GET) gripper, a novel 1-DoF, 3-finger design for securely grasping objects of many shapes and sizes. Mounted on a standard parallel jaw actuator, the design features three narrow, tapered fingers arranged in a two-against-one configuration, where the two fingers converge into a V-shape. The GET gripper is more capable of conforming to object geometries and forming secure grasps than traditional designs with two flat fingers. Inspired by the principle of self-similarity, these V-shaped fingers enable secure grasping across a wide range of object sizes. Further to this end, fingers are parametrically designed for convenient resizing and interchangeability across robotic embodiments with a parallel jaw gripper. Additionally, we incorporate a rigid fingernail for ease in manipulating small objects. Tactile sensing can be integrated into the standalone finger via an externally-mounted camera. A neural network was trained to estimate normal force from tactile images with an average validation error of 1.3 N across a diverse set of geometries. In grasping 15 objects and performing 3 tasks via teleoperation, the GET fingers consistently outperformed standard flat fingers. All finger designs, compatible with multiple robotic embodiments, both incorporating and lacking tactile sensing, are available on GitHub.
Figure 1: Grasp EveryThing (GET) Gripper. (a) Holding a toy hammer with simultaneous tactile image recorded from HDR camera at the back of the single finger side. (b) The GET gripper is designed to be easily integrated onto parallel grippers commonly found on existing in-the-wild data collection systems with added torque constraint. We demonstrate its capability in completing a variety of household tasks through teleoperation on the ALOHA system [1].
Figure 2: Assembly of GET components. Left: Rendered exploded view of mechanical and electronic components in each finger, arranged in order of assembly. Parts that are made of Onyx are 3D printed on a Markforged printer. Right: Fabricated set of GET fingers with rigid fingernails highlighted.
@@ I Introduction
By introducing a third finger, we can create a more robust and universally capable gripper that remains as easy to implement and control. The contact patch provided by this additional finger works to form a lever arm to better resist torques about both normal and frictional directions. With V-shaped fingers, the length of this lever varies as needed in correspondence to the shape or size of grasped objects.
@@ II-A Robotic Gripper Design
Rigid, parallel jaw grippers are easy to model and control due to their simple geometry and 1-DoF. Recently, they have become especially fundamental in the rise of learning-from-demonstration style control strategies that heavily rely on teleoperation and in-the-wild data collection because they are natural to teleoperate [5, 2]. Despite their convenience, traditional flat-fingered parallel jaw grippers often form only two contact patches on grasped objects. This limits their ability to constrain against torques and achieve secure grasps, making them less effective for dynamic tasks and tool usage [7, 10]. Alternative designs have been developed in attempt to overcome this obstacle. Most notable are compliant structure fingers that deform to increase contact area and further geometrically constrain grasped objects [11, 12, 13]. These grippers are favorable due to their improved performance over rigid fingers, but still fail in dynamic tasks requiring strong torque resistance. Other form factors have been developed to increase torque resistance beyond the traditional 2-fingered parallel gripper [1, 14].
By introducing additional DoF, more dexterous, anthropomorphic, and dynamically-capable grippers can be designed [15, 16, 17, 18]. The increased complexity of these grippers empowers more performative action, including strong multi-contact grips and in-hand manipulation behaviors. Yet, hand designs can become bulky, costly, and scale to large numbers of DoF that leave them difficult to control. Crucially, multi-DoF hands are complicated to teleoperate [19, 20], which hinders the development of learning-from-demonstration control policies. Naturally, many have sought to minimize DoFs while maximizing dexterity through underactuation [21]. However, modeling underactuation can be challenging and teleoperation remains nontrivial.
@@ II-B Camera-Based Tactile Sensors
Recently, researchers have developed approaches where a single camera is used to capture both tactile and proximity data, capturing objects and their associated contact geometry in the same image [28, 29]. This eliminates the need for a wrist-mounted camera. However, an externally-mounted camera may be susceptible to lighting changes, introducing noise that could inhibit the ability to reconstruct surface depth as precisely as GelSight sensors traditionally do [23]. This is likely acceptable in learning-based implementations that only utilize RGB tactile images as input, not 3D depth [27, 9, 28].
@@ III-A1 Secure Grasping
Even when grasping more regularly-shaped objects, traditional parallel jaw grippers with two flat fingers struggle to strongly resist torques. We analyze the forces involved in grasping a hammer with a flat, cylindrical handle in Fig. 3b. Here, we assume fingers are not fully rigid and will form patch contacts with the hammer. The 2-finger gripper is able to withstand torque disturbances about the $\hat{x}$ and $\hat{z}$-axes in proportion to the width of the fingers $w$. When a third finger is introduced, the ability to withstand torque additionally depends on the distance between fingers $L$. By designing a parallel gripper with three fingers, we can achieve greater torque compensation about both $\hat{x}$ and $\hat{z}$-axes without the need for overly wide, flat fingers, which may hinder performance in smaller-scale tasks. Resistance to torques about the $\hat{y}$ direction will remain equivalent for both finger configurations and depend on object width.
With our 3-fingered configuration, it is important to note that the greatest resistance against torque disturbances functions about the $\hat{x}$ axis along the length of the fingers. Resistance against disturbances about the grasping axis $\hat{z}$ will rely on friction, with maximum allowed torque scaling down by the frictional coefficient $\mu$. Objects should be grasped strategically to ensure favorable torque compensation.
Beyond incorporating a third finger to improve geometric conformity and increase torque resistance, we design fingers with soft silicone gel pads to further enhance these capabilities. When these soft fingerpads come in contact with a grasped object, they will deform to the local geometry of the object. This works to increase the area of contact patches and grants flexibility in grasping irregularly-shaped objects. Additionally, depending on the geometry of the object, an indentation formed at the contact patch may provide slip resistance along the face of the finger, with the elastic silicone acting as a spring to resist object movement.
@@ III-A2 Self-Similarity & Scalability
On the 2-finger side of our gripper, the fingers come together to form a V-shape. This creates a varying distance $L$ between fingers along their length to enable secure grasping of many object sizes, as shown in Fig. 4. Larger $L$ at the base provides greater compensation against torque disturbances, while smaller $L$ at the tip enables grasping of small objects. With this design, it is expected that larger objects are grasped at the base of the fingers and smaller objects at the tip.
@@
To prevent fingers from bending outward and diverging, which could lead to grasp instability, we pitch them inward by angle $\theta$. Importantly, fingers are designed to interdigitate. They can bend flat to a $0^{\circ}$ pitch angle through contact without interfering with each other, allowing even infinitesimally thin objects to be grasped along the length of the finger.
Fingers are parametrically designed such that they can be resized for compatability with any robotic arm. This is done by adjusting finger length dimension and attachment geometry. Throughout this paper, we refer specifically to fingers designed for the ALOHA system [1]. But, to demonstrate scalability, additional desgins were fabricated for use with a Franka Emika Panda arm, as shown in Fig. 6, and the UMI gripper, as available on GitHub [5].
@@ III-B Optical Design
Camera-based tactile sensing is integrated into the single finger of the gripper. Soldered arrays of RGB LED are placed along the edges of the finger to illuminate the sensor pad. Our finger is significantly longer than it is wide. In compensation, lensed blue LEDs are used at the base of the finger to direct light further along its length. Otherwise, red and green SMD LEDs are placed on either side of the finger and equipped with diffusion filters to promote uniform illumination.
As discussed in Section III-A2, our design is intended to be parametrically scalable. In accordance with this, each LED board could be replaced by flexible COB LED strips, which can be customized to different lengths, eliminating the need to design a new PCB board for each new finger size.
@@ IV-A Tactile Sensing
Tactile images were collected alongside corresponding measurements from a force probe over varying indentation geometries. Indentation geometries include 3 simple shapes and 15 common objects, such as a coin, USB stick, Lego brick, and more. The force probe used for this experiment was designed with an embedded MMS101 6-axis force torque sensor. CAD files for this probe are available on GitHub. To collect data, we pressed our single finger tactile sensor into the probe multiple times for each geometry while streaming tactile images and recording force measurements for each frame, as shown in Fig. 7b. Pressing forces ranged from 0 to 30 N over the collected dataset. In total, 17.5k images were collected. For reference, while streaming at 30 Hz, all of this data can be collected in under 10 min.
@@ IV-B Performance in Grasping Objects
For these experiments, we utilized the ALOHA system [1] retrofitted with an enhanced gripper design from the ALOHA 2 [6]. We compared three different finger configurations: our proposed design, ALOHA ViperX fingers, and a custom-designed set emulating traditional parallel jaw fingers. The traditional fingers are flat and rectangular with a rigid backing and a soft curved gel pad that mimics GET fingers. The ALOHA ViperX fingers are fully rigid with a triangular design that can provide up to four points of contact. We applied high-friction tape across the entire surface of ALOHA ViperX fingers to maximize their capabilities.
@@ IV-C Performance in Manipulation Tasks
Task performance for each set of fingers is plotted in Fig. 10. Our fingers outperformed baseline fingers, specifically in the Toy Hammer and Peanut Butter Spread tasks. Traditional fingers struggled to resist torque moments in these tasks given only two small contact patches. Furthermore, they could not reliably pick small objects in the Car Part Assembly task. ALOHA ViperX fingers performed better in the Car Part Assembly task, but struggled in the Peanut Butter Spread task. These rigid fingers could not conform to the butter knife’s convex geometry to securely grasp it, as displayed in Fig. 3a.
The GET fingers more successfully completed all tasks than baseline fingers. In the Car Part Assembly task, their fingernails helped pick small pieces from clutter and place them firmly into the desired location. In the Peanut Butter Spread task, the GET design could securely grasp the irregularly-shaped butter knife through contact with the cracker. Fingernails were also helpful in this task for removing the peanut butter lid and picking crackers from the stack. In the Toy Hammer task, GET fingers outperformed the ALOHA ViperX fingers, even with the hammer’s flat, cylindrical geometry. GET fingers had greater capacity to resist high torque required to pry the nail from the board, as shown in Fig. 3b. In all tasks, we observe decreased completion time with the GET fingers, suggesting greater teleoperator confidence and ease of use.
@@ V Conclusion
We present the GET gripper, a novel 3-fingered design for parallel jaw actuators to more securely grasp objects of many shapes and sizes. This design has been shown to outperform standard 2-fingered designs, particularly in grasping small objects and securely holding larger ones, such as tools for dynamic tasks. Additionally, high-resolution tactile imaging is integrated into fingers with force sensing capabilities. The GET gripper could be deployed on teleoperation interfaces to collect rich data for training policies over a wide range of tasks. Designs for use with ALOHA, Franka Emika Panda, and UMI systems are available on GitHub, including variants without tactile sensing or silicone gels for rapid manufacturing.